Test Ho: µ =100; H1: µ < 100, using n = 36 and alpha = .05 If the sample mean=92 and the sample standard deviation = 18, which of the following is true?

  • A. test statistic = -2.67; critical value = 1.69; we fail to reject Ho.

  • B. test statistic = -2.67; critical value = 1.96; we fail to reject Ho.

  • C. test statistic = -2.67; critical value = -1.96; we reject Ho.

  • D. test statistic = -2.67; critical value = -1.69; we reject Ho.
Answer #1

answer....

Consider the given problem here the null hypothesis is given below.

=> H0:µ = 100 and the corresponding alternative hypothesis is H1: µ < 100.

So, the test statistic is given by, “t” = √n*[mean(x) – µ ] / sd = 6*(92-100)/18 = (-2.67).

Here the alternative hypothesis is “H1: µ < 100”, => it’s a left tail test, => at the given level of significance (5%) the critical value is “-1.69” (use “t” table here level of significance is 5% and the degree of freedom is 35).

=> “D” be the correct option.